OpenStudy (anonymous):

ALGEBRA QUICK QUESTION: how do i know which order to do things when solving a one-variable equation?? example of a problem: 7x - (2x - 5) = 12x - 3 - 3x

OpenStudy (amistre64):

thre is no specific order you have to approach it from

OpenStudy (amistre64):

as a rule of thumb tho, its easier to play with things if they are all out in the open and not hiding inside of groups

OpenStudy (anonymous):

oh okay!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thanks but i think it does matter because i just did it and i got 1.6 instead of the right answer (2)

OpenStudy (amistre64):

show me what you did

OpenStudy (amistre64):

its not that a specific order matters, as long as you do the proper maths ....

OpenStudy (amistre64):

7x - (2x - 5) = 12x - 3 - 3x 7x - (2x - 5) = 12x - 3 - 3x + 7x - 7x 7x - (2x - 5) = 7x + 12x - 3 -10x + 5 - 5 7x - (2x - 5) = 7x + 5 + 2x - 8 +2x - 2x 7x - (2x - 5) = 7x -2x + 5 + 4x - 8 7x - (2x - 5) = 7x -(2x - 5) + 4x - 8 well, this is only true of 4x-8 = 0

OpenStudy (amistre64):

sis i follow any 'proper' order? or did i just do proper maths?

OpenStudy (amistre64):

*did i follow

OpenStudy (anonymous):

Sorry i haven't replied faster but I get confused when its solved like that and it confused me even more with that question.. so just nevermind thanks for trying to help me!!

OpenStudy (amistre64):

i was simply pointing out that there is no 'specific' way to approach the solution. you simply work proper maths. a good process would be: distribute on the left collect all like terms to one side and the other combine them then divide off the excess and you are left with x

OpenStudy (amistre64):

7x - (2x - 5) = 12x - 3 - 3x 7x - 2x -- 5 = 12x - 3 - 3x 7x - 2x + 5 = 12x - 3 - 3x 7x - 2x -12x +3x = - 3 - 5 5x -12x +3x = - 8 8x -12x = - 8 -4x = - 8 x = -8/-4